Suppose that monopolist has a demand function p1= 100-Q1 , p1=80-Q1 , C=6Q1+6Q2. How much should be sold in the market? What are corresponding prices? Find total profit?


1
Expert's answer
2021-06-02T11:10:25-0400
"MR_1=100-Q_1,"

"MR_2=80-Q_2,"

Set "MR_1=MR_2=MC"


"C=6Q_1+6Q_2=6(Q_1+Q_2)"

"Q_1+Q_2=Q"

"C=6Q"

"MC=6"


"MR_1=6=>100-Q_1=6"

"Q_1=94"

"p_1=100-94=6"



"MR_2=6=>80-Q_2=6"

"Q_2=74"

"p_2=80-74=6"

Profits are the difference between the sum of the revenues generated in both market segments and the total cost of producing the output:


"\pi=p_1(MR_1)+p_2(MR_2)-C"



"\pi=6(6)+6(6)-6(6+6)=0"